OpenStudy (anonymous):

A positive number is divided by 5 instead of being multiplied by 5. what % is the result of the required correct value.

OpenStudy (anonymous):

good medels as reward :)

myininaya (myininaya):

can you please restate the question (second sentence)? i don't understand it

OpenStudy (stacey):

I think you mean: What percent of the correct value is the result? correct value = 5x result = x/5 Percent * correct = result Percent = result/correct = (x/5)/5x

myininaya (myininaya):

x/5 divded by 5x =x/5 * 1/(5x) =x/(5*5x)=x/(25x)=1/25=4%

OpenStudy (anonymous):

yes its 4% :)